# Strategic Client Selection to Address Non-IIDness in HAPS-enabled FL Networks

### v2

@@ Abstract
The deployment of federated learning (FL) in non-terrestrial networks (NTN) that are supported by high-altitude platform stations (HAPS) offers numerous advantages. Due to its large footprint, it facilitates interaction with a large number of line-of-sight (LoS) ground clients, each possessing diverse datasets along with distinct communication and computational capabilities. The presence of many clients enhances the accuracy of the FL model and speeds up convergence. However, the variety of datasets among these clients poses a significant challenge, as it leads to pervasive non-independent and identically distributed (non-IID) data. The data non-IIDness results in markedly reduced training accuracy and slower convergence rates. To address this issue, we propose a novel weighted attribute-based client selection strategy that leverages multiple user-specific attributes, including historical traffic patterns, instantaneous channel conditions, computational capabilities, and previous-round learning performance. By combining these attributes into a composite score for each user at every FL round and selecting users with higher scores as FL clients, the framework ensures more uniform and representative data distributions, effectively mitigating the adverse effects of non-IID data. Simulation results corroborate the effectiveness of the proposed client selection strategy in enhancing FL model accuracy and convergence rate, as well as reducing training loss, by effectively addressing the critical challenge of data non-IIDness in large-scale FL system implementations.
@@ I Introduction
In the rapidly evolving domain of distributed machine learning, federated learning (FL) has emerged as a paradigm-shifting approach, particularly suited for large-scale systems [1]. Characterized by its collaborative yet decentralized nature, FL develops a global learning model using the multitude of data repositories located far apart while preserving their privacy [2]. The integration of aerial devices from non-terrestrial networks (NTN), such as high altitude platform station (HAPS), into FL systems, is increasingly recognized as essential [3][4]. With its extensive coverage area and dominant line-of-sight (LoS) links, HAPS provides broader geographical reach and lower latency compared to satellites [5]. This capability enables a wide array of ground devices to serve as FL clients, thereby fostering a robust, large-scale FL network [6].
@@
Expanding the FL to a large-scale system yields substantial benefits, including enhanced model accuracy and faster convergence [7][8]. However, this scalability also introduces the challenge of non-independently and identically distributed (non-IID) data. This issue becomes even more severe in highly heterogeneous NTN, where users exhibit significant variations in behavior, data sources, and usage patterns [9]. Non-IID data basically arises when each client’s dataset differs substantially from that of other clients, rather than originating from a shared distribution. These differences are often driven by unique usage behaviors, contextual factors, or sensed environmental data, resulting in substantial variability in data distributions across users. Consequently, selecting these users with non-IID data distributions to participate in the FL training process degrades the overall effectiveness of the learning, resulting in slower convergence rates and reduced model accuracy [10].
@@
To address this challenge, some recent works [10, 12, 11, 13, 14, 15] have proposed a variety of client selection strategies in FL systems, including clustering-based approaches, reputation-based frameworks, game-theoretic methods, stochastic, and energy-efficient algorithms. These strategies primarily optimize FL systems for operational constraints such as efficiency and reliability. Among these strategies, cluster FL has emerged as a promising solution for addressing non-IID data by grouping clients with similar data distributions [10][11]. For instance, [15] leverages advanced K-means clustering to group clients based on feature similarity, while [11] uses hierarchical clustering for the same purpose. However, these approaches generally focus on intrinsic dataset features, which may be less accessible to an FL server in practice, overlooking user network behavior and traffic characteristics.
@@
Unlike previous approaches, we propose a weighted attribute-based strategy to select clients with more homogeneous data distributions (i.e., less non-IID). Within this strategy, the FL server leverages several user-specific attributes, including historical traffic patterns, instantaneous channel conditions, computational capabilities, and previous-round learning performance during FL training to assign a composite score to each user at every FL round. Users who achieve higher composite scores tend to have similar data distributions, making them prime candidates for FL training.
@@
The rationale for selecting these attributes offers two key benefits. First, it enables the identification of users with similar datasets by analyzing historical traffic patterns that reflect general data usage and user types. This approach provides insights into the potential datasets and previous learning performance, highlighting each user’s prior contributions to the global model. Second, it focuses on identifying the strongest users among those with similar datasets by examining their instantaneous channel conditions, which indicate their communication reliability. Additionally, considering their computational capabilities ensures efficient local training and helps mitigate the effects of stragglers. By combining these attributes into a single weighted composite score, the server can effectively pinpoint the most capable clients whose data is more uniform or similar. This reduces the non-IIDness of the datasets and enhances the performance of FL training.
As the central node, the FL server generally collects comprehensive data on user behavior and network traffic [16], including daily usage patterns and preferred applications. It also estimates instantaneous channel state information (CSI) before any transmission and is notified of users’ computational capabilities (e.g., CPU speed/frequency [7]). Moreover, the server maintains a record of each user’s learning performance during every round of the FL training process.
@@ II-A Network Topology Model
We consider an NTN architecture, as shown in Fig. 1, which consists of a HAPS located in the stratosphere and $K$ ground users [^1] Throughout this work, we use the term “client” to refer to any ground user that has been selected to participate in FL training. that can participate in the FL process. The HAPS acts as an FL server, offering extensive coverage and facilitating communication with the geographically distributed ground users within its footprint. We assume that the locations of these ground users follow a uniform distribution. Each user has local data and is equipped with computational resources, enabling them to undertake local training as part of the FL framework. It is assumed that the HAPS is equipped with high computational capabilities and has access to historical network data, such as user traffic behavior and communication/computation capability metrics, through integrated ground stations or satellite links. These features make HAPS well-suited for large-scale FL deployments, providing a centralized node for client selection and global model aggregation.
@@ II-B Network Traffic Features Modeling
$\lambda_{k}\;=\;\lambda_{max,k}\,\bigl{(}1-P_{\text{loss},k}\bigr{)}.$ (4)
Here, $P_{\text{loss},k}$ is the probability that a packet is lost due to bit errors. For a packet of size $S_{k}$ bits, and assuming independence of bit errors, we can approximate
@@
Under a Rician fading model, $|h_{k}|^{2}$ follows a noncentral chi-square distribution (with 2 degrees of freedom) that combines a deterministic LoS component and a stochastic non-LoS (NLoS) component [19]. In a high-SNR regime, the deterministic LoS component typically dominates; thus, one can approximate $|h_{k}|^{2}$ by its mean $\Omega_{k}$, where
@@ III-A Selection Criteria
Traffic History Metrics
We focus on two indicators, including the traffic volume $V_{k}$ and burstiness $B_{k}$. Larger $V_{k}$ typically means higher traffic rates, and higher $B_{k}$ signifies more bursty (less stable) traffic. For client $k$, we combine these into a single metric $\hat{t}$ by first normalizing both $V_{k}$ and $B_{k}$ to $[0,1]$, then defining
@@
We represent the channel quality of each client by a normalized measure $\hat{r}_{k}\in[0,1]$, which can be infer from the signal-to-noise ratio as a function of channel gain, i.e., $\frac{p_{k}|h_{k}|^{2}}{N_{0}b_{k}}$. A larger $\hat{r}_{k}$ indicates better channel quality and higher potential throughput. To account for this attribute, $\hat{r}_{k}$ is incorporated into the final composite score to ensure that users with better wireless channel quality are more likely to be selected as FL clients.
@@
We consider client $k$ with CPU computing capability $f_{k}$ (in CPU cycles per second) and let $C_{k}$ denote the number of CPU cycles required to process one data sample. At any communication round, for a given local computation time $l_{k}$ and $J_{k}$ data samples, $f_{k}$ is determined by
@@
FL Performance Score
We track each client’s previous-round FL learning performance via a dynamic improvement metric $\hat{m}_{k}$. Following local training in round $n$, each selected client $k$ contributes to the FL training process with a normalized improvement $\Delta_{k}^{(n)}\in[0,1]$, which is defined as a local loss reduction. We keep a running dynamic performance score $\hat{m}_{k}$ for each round as
$\hat{m}_{k}^{(n+1)}\;\leftarrow\;\zeta\,\hat{m}_{k}^{(n)}\;+\;\bigl{(}1-\zeta\bigr{)}\,\Delta_{k}^{(n)},\quad 0<\zeta<1,\>$ (14)
where $\zeta$ is a tunable scaler parameter. Larger $\zeta$ emphasizes past values of $m_{k}$, and smaller $\zeta$ highlights the latest improvements. Therefore, clients exhibiting repeatedly high improvements maintain a higher $m_{k}$.
@@ III-B Composite Score Formulation
After each attribute is normalized to $\hat{t}_{k}$, $\hat{r}_{k}$, $\hat{f}_{k}$, and $\hat{m}_{k}$, we compute weighted attribute composite score as follows:
$\mathrm{score}_{k}\;=\;\varepsilon_{t}\,\bigl{(}1-\hat{t}_{k}\bigr{)}\;+\;\varepsilon_{r}\;\hat{r}_{k}\;+\;\varepsilon_{f}\;\hat{f}_{k}\;+\;\varepsilon_{m}\;\hat{m}_{k},\quad\forall\,k,$ (15)
where $\varepsilon_{(.)}$ are weights that are assigned to each criteria. For tractability, we assume equal weights for all criteria, i.e., $\varepsilon_{t}=\varepsilon_{d}=\varepsilon_{r}=\varepsilon_{f}=\varepsilon_{m}=\varepsilon$. In each round, we select a subset $\mathcal{S}$ of users with the highest composite scores to serve as FL clients. This subset is formally defined as
$\mathcal{S}=\{k|\forall k,\>\mathrm{score}_{k}\geq\mathrm{score}_{\text{th}}\},$ (16)
where $\text{score}_{\text{th}}$ is a predefined selection threshold. By prioritizing users with higher scores, the selection process favors clients whose datasets tend to exhibit greater similarity, thereby reducing data heterogeneity and mitigating the impact of non-IID distributions across the selected participants.
@@
1:  Initialization: Total number of users $K$, composite score weights $(\varepsilon_{t},\,\varepsilon_{r},\,\varepsilon_{f},\,\varepsilon_{m})$, threshold $\text{score}_{\text{th}}$.
@@
3:     Calculate the attribute associated with each selection criterion: $(\hat{t}_{k},\hat{r}_{k},\hat{f}_{k},\hat{m}_{k})$
4:     Compute composite score for each user using eq. (15).
5:  end for
6:  Select the users with the highest score as FL clients:$\mathcal{S}\;\leftarrow\;\{k|\forall k,\>\text{score}_{k}\geq\text{score}_{\text{th}}\}.$
7:  Return Subset $\mathcal{S}$.
@@ IV Augmented FL Model
    where $\tau$ is the FedProx parameter and $\mathcal{L}_{\mathrm{CE}}$ is the local loss (e.g., cross-entropy). This yields an updated local model $\mathbf{w}_{k}^{(n)}$.
@@
1:  //  Initialization: Training set $\mathcal{D}$, test set $\mathcal{T}$, total clients $K$, total FL rounds $N$, local epochs $E$, learning rate $\eta$, batch size $B$, FedProx parameter $\tau$, constant $\zeta$.
@@
14:        Obtain local model $\mathbf{w}_{k}^{(n)}$, local loss $L_{k}^{(n)}$, and improvement metric $\Delta_{k}^{(n)}$.
@@ V Simulation Results
In this section, we evaluate the proposed weighted attribute-based client selection strategy by employing the augmented FL system in a HAPS-aided NTN. The setup includes $500$ ground users ($K=500$), each with non-IID data distributions to reflect realistic usage patterns. In particular, we use the non-IID CIFAR-10 dataset to assess FL training performance under simulated user traffic behaviors. To emulate real-world conditions, synthetic features such as burstiness and traffic volume are derived from modeled Rician fading channels and log-normal packet size distributions, capturing key variations in user behavior. Our primary focus is to investigate how the client selection strategy influences FL outcomes. TABLE I lists the simulation parameters and their values. For a fair comparison, we evaluate our approach against two widely used FL client selection strategies: (i) Random Strategy, which serves as the baseline benchmark commonly adopted in terrestrial FL systems, and (ii) Resource-Aware Strategy, which clusters and selects clients based on their available communication resources (e.g., transmission power, bandwidth) and computational resources (e.g., CPU speed) [22]. All figures in this section present a comparative analysis of our proposed strategy against the two benchmark approaches.
In Fig. 2, we examine the average FL test accuracy over communication rounds. As the number of rounds increases, the proposed client selection strategy demonstrates a notable improvement in accuracy, diverging significantly from the benchmark trends. The two benchmark strategies show comparable performance, with the Resource-Aware Strategy offering a slight edge over the Random Strategy. After $n=5$ rounds, our approach outperforms the benchmarks by up to $34\%$, underscoring its effectiveness and high performance.
@@
Figure 4: Average FL convergence rate over communication rounds.
VI Conclusion

### v3

@@ Abstract
Federated learning (FL) supported by high-altitude platform stations (HAPS) in non-terrestrial networks (NTN) provides a distinct advantage by allowing the participation of a large number of line-of-sight (LoS) ground clients compared to terrestrial networks, due to their extensive coverage. The presence of many clients enhances the accuracy of the FL model and speeds up convergence. However, the diverse datasets among these clients result in non-independently and identically distributed (non-IID) data, which can negatively affect training accuracy and lead to slower convergence rates. To address this issue, we propose a novel weighted attribute-based client selection strategy that leverages multiple user-specific attributes, including historical traffic patterns, channel qualities, computational capabilities, and prior-round learning performance. By combining these attributes into a composite score for each user at every FL round and selecting users with higher scores as FL clients, the framework ensures more uniform and representative data distributions. Simulation results corroborate the effectiveness of the proposed client selection strategy in enhancing FL model accuracy and convergence rate, as well as reducing training loss.
@@ I Introduction
In the rapidly evolving domain of privacy-preserving machine learning, federated learning (FL) has emerged as a paradigm-shifting approach that develops a global learning model by collaborating among the multitude of clients with unshared data repositories  [1, 2]. The larger the number of repositories, the better FL performs. To achieve this performance, integrating aerial devices from non-terrestrial networks (NTN), such as high altitude platform stations (HAPS), into FL systems has been suggested [3][4]. Because HAPS is located in the stratosphere, it offers a broader geographical reach and direct line-of-sight (LoS) links with ground devices [5]. This capability substantially broadens the set of eligible FL clients by enabling a diverse range of ground devices to participate—an advantage that terrestrial networks typically cannot offer due to their limited coverage and LoS links [6].
@@
However, the large number of clients also introduces the challenge of non-independently and identically distributed (non-IID) data, which arises when each client’s dataset differs substantially from that of other clients, rather than originating from a shared distribution [7]. These differences are often driven by unique traffic usage behaviors, contextual factors, or sensed environmental data, resulting in substantial variability in data distributions across users. Selecting the users with non-IID data distributions to participate in the FL training process degrades the overall effectiveness of the learning, resulting in slower convergence rates and reduced model accuracy [8].
@@
To address this challenge, some recent works [8, 10, 9, 11, 12, 13] have proposed a variety of client selection strategies for FL. Focusing primarily on optimizing operational constraints like efficiency and reliability for FL, these strategies incorporate methods such as client clustering, reputation, game-theoretic, and stochastic techniques. Among these, cluster FL has emerged as a promising solution for addressing non-IID data by grouping clients with similar data distributions [8][9]. For instance, [13] leverages advanced K-means clustering to group clients based on feature similarity, while [9] uses hierarchical clustering for the same purpose. These approaches primarily focus on features derived from the user’s local raw dataset, such as gradients or model updates. However, they often overlook important aspects of the user’s historical network traffic behavior and characteristics, such as past traffic volume and burstiness, which are strongly correlated with their local datasets [14][15].
Figure 1: Network model.
@@
In this work, we present our primary contribution: a weighted attribute-based strategy that selects a more homogeneous set of users by considering features derived from both the user’s local dataset and historical network traffic. This strategy computes a composite score for each user in every communication round by leveraging user-specific attributes, including historical traffic patterns, channel qualities, computational capabilities, and prior-round learning performance. Users with higher composite scores are more likely to possess similar data distributions, making them ideal participants in FL training.
@@
The rationale for selecting these attributes is twofold. First, they help identify users with similar datasets by analyzing historical traffic patterns. This analysis provides insights into traffic usage and users’ prior behavior on the network. Second, the selection focuses on the strongest users among those with similar datasets by assessing their current channel conditions, computational capabilities, and previous-round learning performance, ensuring efficient local training while reducing the impact of stragglers. By consolidating these attributes into a single weighted score, the server can effectively identify the most capable clients with more uniform data, thereby enhancing the performance of FL training.
@@ II-A Network Topology Model
We consider an NTN architecture, as shown in Fig. 1, which consists of a HAPS located in the stratosphere and $K$ ground users [^1] Throughout this work, we use the term “client” to refer to any ground user that has been selected to participate in FL training. that can participate in the FL process. The HAPS acts as an FL server, offering extensive coverage and facilitating communication with the geographically distributed ground users within its footprint. We assume that the locations of these ground users follow a uniform distribution. Each user has local data and is equipped with computational resources, enabling them to undertake local training as part of the FL framework. We assume that HAPS functions as an aerial base station within the proposed NTN architecture, thus having access to historical user network behavior, including user traffic and service demand in terms of traffic volume and burstiness.
@@ II-B Network Traffic Features Modeling
$\lambda_{k}\;=\;\lambda_{max,k}\,\bigl{(}1-P_{\text{loss},k}\bigr{)},$ (4)
where $P_{\text{loss},k}$ is the probability that a packet is lost due to bit errors. For a packet of size $S_{k}$ bits, and assuming independence of bit errors, we can approximate
@@
Under a Rician fading model, the dominant LoS HAPS links cause $|h_{k}|^{2}$ to follow a noncentral chi-square distribution with 2 degrees of freedom, which combines a deterministic LoS component and a stochastic non-LoS (NLoS) component [17]. In a high signal-to-noise ratio (SNR) regime, the deterministic LoS component typically dominates; thus, one can approximate $|h_{k}|^{2}$ by its mean $\Omega_{k}$, where
@@ III-A Selection Criteria
Historical Traffic Pattern
To accurately capture users’ historical traffic behavior, we focus on two key metrics: traffic volume $V_{k}$ and burstiness $B_{k}$, both of which can be derived from historical network traffic records available at HAPS. Larger $V_{k}$ typically means higher traffic rates, and higher $B_{k}$ signifies more bursty (less stable) traffic. For client $k$, we combine these into a single metric $\hat{t}$ by first normalizing both $V_{k}$ and $B_{k}$ to $[0,1]$, then defining
@@
During the FL process, the channel quality for each client is represented by a normalized measure $\hat{r}_{k}\in[0,1]$, which can be inferred from the available SNR as a function of channel gain at HAPS, i.e., $\frac{p_{k}|h_{k}|^{2}}{N_{0}b_{k}}$. A larger $\hat{r}_{k}$ indicates better channel quality and higher potential throughput. To account for this attribute, $\hat{r}_{k}$ is incorporated into the final composite score to ensure that users with better wireless channel quality are more likely to be selected as FL clients.
@@
During the FL process, the computing capability $f_{k}$ (in CPU cycles per second) of client $k$ can be determined by HAPS [20]. Let $C_{k}$ denote the number of CPU cycles required to process one data sample. At any communication round, for a given local computation time $l_{k}$ and $J_{k}$ data samples, $f_{k}$ is determined by
@@
Prior-round Learning Performance
During the FL process, HAPS can track each client’s previous-round FL learning performance via a dynamic improvement metric $\hat{m}_{k}$ defined in the following Theorem 1.
Theorem 1.
Let $\Delta_{k}^{(n)}\!\in[0,1]$ be the normalized loss reduction reported by client $k$ after local training round $n$ and choose a memory factor $\zeta\in(0,1)$. Define the recursive update
$\hat{m}_{k}^{(n+1)}\;=\;\zeta\,\hat{m}_{k}^{(n)}+(1-\zeta)\,\Delta_{k}^{(n)}.\vspace{-2mm}$ (14)
Then
  1. (i)
    Optimality. $\hat{m}_{k}^{(n+1)}$ is the unique minimiser of the exponentially weighted least-squares cost $J_{n+1}(\theta)=\sum_{t=0}^{n}\zeta^{\,t}(\theta-\Delta_{k}^{(n-t)})^{2}$.
  2. (ii)
    Closed-form. $\hat{m}_{k}^{(n+1)}=(1-\zeta)\sum_{t=0}^{n}\zeta^{\,t}\Delta_{k}^{(n-t)}+\zeta^{\,n+1}\hat{m}_{k}^{(0)}$, i.e., an exponential moving average (EMA) of past improvements with decay $\zeta$.
  3. (iii)
    Boundedness. If $0\leq\hat{m}_{k}^{(0)}\leq 1$, then $0\leq\hat{m}_{k}^{(n)}\leq 1$ for every $n$.
  4. (iv)
    Convergence. If $\{\Delta_{k}^{(n)}\}$ is mean-stationary with mean $\mu^{\prime}_{k}$, then $\mathbb{E}[\hat{m}_{k}^{(n)}]\to\mu^{\prime}_{k}$ as $n\to\infty$.
Proof.
See Appendix.
Intuitively, a smaller $\zeta$ emphasises the most recent round, whereas a larger $\zeta$ rewards sustained performance.
@@ III-B Composite Score Formulation
After calculating and normalizing all the attributes $\hat{t}_{k}$, $\hat{r}_{k}$, $\hat{f}_{k}$, and $\hat{m}_{k}$ so that higher values are always considered better, HAPS computes a weighted composite score for each attribute using the following linear equation:
$\mathrm{score}_{k}=\varepsilon_{t}\hat{t}_{k}+\varepsilon_{r}\hat{r}_{k}+\varepsilon_{f}\hat{f}_{k}+\varepsilon_{m}\hat{m}_{k},\quad\forall k,$ (15)
This score remains within the range of $[0,1]$ due to the properties of convex combinations. Weights $\varepsilon_{(.)}$ are assigned to each criterion. After each communication round, the absolute Pearson correlation $\beta_{i}$ is calculated for each attribute $i\in\{t,r,f,m\}$ in relation to the reduction in global validation loss on the FL server that was achieved during that round. The corresponding decrease in global validation loss is then recorded as follows:
$y^{(n)}=\frac{\mathcal{L}_{\mathrm{val}}^{(n-1)}-\mathcal{L}_{\mathrm{val}}^{(n)}}{\mathcal{L}_{\mathrm{val}}^{(n-1)}},$ (16)
where $\mathcal{L}_{\mathrm{val}}^{(n)}$ is the global validation loss after the $n$-th aggregation.
Given a memory window of the last $H$ rounds, a paired dataset $\{(x^{(n)}_{i},y^{(n)})\}_{n=n_{0}}^{n_{0}+H-1}$ can be created for every attribute $i$. The relevance score for each attribute is calculated using the absolute Pearson correlation $\beta_{i}=\left|\mathrm{corr}\left(\{x^{(n)}_{i}\}_{n=n_{0}}^{n_{0}+H-1},\{y^{(n)}\}_{n=n_{0}}^{n_{0}+H-1}\right)\right|$. Consequently, the weight vector $\bm{\varepsilon}$ is updated online as
$\varepsilon_{i}=\frac{\beta_{i}}{\sum_{j\in\{t,r,f,m\}}\beta_{j}},\quad i\in\{t,r,f,m\}.$ (17)
This procedure, which operates in $O(|\mathcal{K}|)$, adaptively increases the weight of factors that demonstrate higher predictive power regarding the progress of FL training while ensuring that the constraint $\sum_{i}\varepsilon_{i}=1$ is maintained.
As a result, at round $n$, clients with a composite score exceeding a predefined threshold $\mathrm{score}_{\mathrm{th}}$ are selected
$\mathcal{S}^{(n)}=\{k\mid\mathrm{score}_{k}^{(n)}\geq\mathrm{score}_{\mathrm{th}}\}.$ (18)
By prioritizing clients with higher scores, this selection process favors users whose datasets tend to be more similar. It is important to note that while this approach decreases the number of users participating in the HAPS-aided FL process, the number of selected users remains significantly higher than that of the terrestrial counterpart.
@@
1:  Initialization: Total number of users $K$, threshold $\text{score}_{\text{th}}$.
@@
3:     Calculate the attribute associated with each selection criterion: $(\hat{t}_{k},\hat{r}_{k},\hat{f}_{k},\hat{m}_{k})$.
4:     Calculate the associated weight vector $\mathbf{\epsilon}=\{\varepsilon_{t},\,\varepsilon_{r},\,\varepsilon_{f},\,\varepsilon_{m}\}$.
5:     Compute composite score for each user using eq. (15).
6:  end for
7:  Select the users with the highest score as FL clients:$\mathcal{S}\;\leftarrow\;\{k|\text{score}_{k}\geq\text{score}_{\text{th}}\}.$
8:  Return Subset $\mathcal{S}$.
@@ IV Augmented FL Model
    where $\tau$ is a constant parameter and $\mathcal{L}_{\mathrm{CE}}$ is the local loss (e.g., cross-entropy). This yields an updated local model $\mathbf{w}_{k}^{(n)}$.
@@
1:  //  Initialization: Training set $\mathcal{D}$, test set $\mathcal{T}$, total clients $K$, total FL rounds $N$, local epochs $E$, learning rate $\eta$, batch size $B$, constant parameter $\tau$, constant $\zeta$.
@@
14:        Obtain local model $\mathbf{w}_{k}^{(n)}$, local loss $\mathcal{L}_{k}$, and improvement metric $\Delta_{k}^{(n)}$.
@@ V Simulation Results
In this section, we evaluate the proposed weighted attribute-based client selection strategy using the augmented FL system within a HAPS-aided NTN. The setup involves 500 ground users ($K=500$), each with non-IID data distributions to reflect realistic usage patterns. Specifically, for the non-IID split, we employ a standard label-shard procedure: the 50,000 CIFAR-10 training images are sorted by class, divided into 1,000 equal-sized shards, and two random shards are assigned to each of the 500 clients. As a result, each client’s local dataset contains images from at most two classes. This setup creates a strong, widely used label-skew non-IID scenario that is fully reproducible and comparable with prior FL research.
To emulate real-world conditions, we generate synthetic features such as burstiness and traffic volume, modeled using Rician fading channels and log-normal packet size distributions. This approach captures key variations in user historical traffic behavior. Our primary focus is to investigate how the client selection strategy impacts the outcomes of FL. TABLE I lists the simulation parameters and their values. For a fair comparison, we evaluate our approach against two widely used FL client selection strategies: (i) Random Selection Strategy, which serves as the baseline benchmark commonly adopted in terrestrial FL systems and the idea is to select clients randomly regardless of their resource availability or channel condition, and (ii) Resource-Aware Strategy, which clusters and selects clients based on their available communication resources (e.g., transmission power, bandwidth) and computational resources (e.g., CPU speed) [21]. All figures in this section present a comparative analysis of our proposed strategy against the two benchmark approaches.
In Fig. 2, we examine the average FL test accuracy over communication rounds. As the number of rounds increases, the proposed client selection strategy demonstrates a notable improvement in accuracy, diverging significantly from the benchmark trends. The two benchmark strategies show comparable performance, with the Resource-Aware Strategy offering a slight edge over the Random Selection Strategy. After $n=5$ rounds, our approach outperforms the benchmarks by up to $34\%$, underscoring its effectiveness and high performance.
@@
Figure 4: Average FL convergence rate over communication rounds.
VI Conclusion
Proof.
(i) Optimality. Define the normalised, exponentially–weighted least-squares cost
$J_{n+1}(\theta)\;=\;\sum_{t=0}^{n}(1-\zeta)\,\zeta^{t}(\theta-\Delta_{k}^{(n-t)})^{2},\qquad 0<\zeta<1,$ (21)
so that the weights sum to one. Setting $\partial J_{n+1}/\partial\theta=0$ yields $\theta^{\star}=(1-\zeta)\sum_{t=0}^{n}\zeta^{t}\Delta_{k}^{(n-t)},$ which is exactly the EMA update $\hat{m}_{k}^{(n+1)}=\zeta\hat{m}_{k}^{(n)}+(1-\zeta)\Delta_{k}^{(n)}.$
(ii) Closed-form expansion. Substituting $\hat{m}_{k}^{(n)}$ recursively, $\hat{m}_{k}^{(n+1)}=\zeta\bigl{[}\zeta\hat{m}_{k}^{(n-1)}+(1-\zeta)\Delta_{k}^{(n-1)}\bigr{]}+(1-\zeta)\Delta_{k}^{(n)}$ and iterating to $\hat{m}_{k}^{(0)}$ gives
$\hat{m}_{k}^{(n+1)}=(1-\zeta)\sum_{t=0}^{n}\zeta^{t}\Delta_{k}^{(n-t)}+\zeta^{\,n+1}\hat{m}_{k}^{(0)},$ (22)
which is the standard finite-horizon EMA form.
(iii) Boundedness. Assuming the initial value obeys $0\leq\hat{m}_{k}^{(0)}\leq 1$, all weights in the EMA above are non-negative and sum to 1, while every $\Delta_{k}^{(t)}\in[0,1]$. Hence $0\leq\hat{m}_{k}^{(n)}\leq 1$ for every $n$, i.e., the score remains bounded and interpretable.
(iv) Convergence. Let $E_{n}=\mathbb{E}[\hat{m}_{k}^{(n)}]$ and assume the improvement sequence $\{\Delta_{k}^{(n)}\}$ is mean-stationary with $\mu^{\prime}_{k}=\mathbb{E}[\Delta_{k}^{(n)}]$. Taking expectations in the update gives the linear recursion $E_{n+1}=\zeta E_{n}+(1-\zeta)\mu^{\prime}_{k}.$ Its solution is
$E_{n}=\mu^{\prime}_{k}\bigl{(}1-\zeta^{\,n}\bigr{)}+\zeta^{\,n}\hat{m}_{k}^{(0)}\;\xrightarrow[n\to\infty]{}\;\mu^{\prime}_{k},$ (23)
because $10<\zeta<$ 1. Thus, the expected score converges to the long-run mean improvement $\mu^{\prime}_{k}$. ∎